OpenStudy (anonymous):

Ok im stuck! Nature of the roots of x^2=0 a: no real roots c: one real roots b: two distinct real roots d: three distinct real roots

OpenStudy (amistre64):

what are the solutions to x^2 = 0?

OpenStudy (solomonzelman):

\[x^2=0.\] you know that when you have an equation like\[x^2=c~~~~~then,~~~~x=±\sqrt{c}\]So in this cae it ould be just zero. Because there is no such a thing as +0 or -0.

OpenStudy (mathstudent55):

Here's the good, old quadratic formula. \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \) The part \(b^2 - 4ac\) is called the discriminant, \(d\). It allows you to tell the nature of the roots of the quadratic equation without having to solve the quadratic equation. To do so, first, evaluate the discriminant, \(d\). Then if: \(d \lt 0\) you get two complex roots. \(d = 0\) you get one single real root. \(d \gt 0\) you get two real roots.

OpenStudy (mathstudent55):

In your case, a = 1, b = 0, and c = 0, so d = 0.